please explain hyperconjucation

Grade:11

2 Answers

sarthak mitra
12 years ago

hyperconjugation is a special type of resonance (no bond resonance)here the H atom carries the +ve charge but there is no bond between C & H.there should be at least 1 H atom in alpha C w.r.t( C=C )                                                                                                                    

                                                     C-C-CH=CH2

                                                     1 2  3

2nd C is alpha C . as there is no H there no hyperconjugation occurs.

 

                                                    C-CH2-CH=CH2    after Hc  C-CH-CH=CH2  

                                                                                              H(+)

                                                    1  2     3

at alpha position (2nd C wrt C=C) hyperconjugable H is present so hyperconjugation occurs. 

no.of hyper conjugable  H present in a compound sometimes helps in detecting the stability of the compound.

Godfrey Classic Prince
12 years ago

Dear moinack shaikh,

In organic che,mistry, hyperconjugation is the interaction of the electrons in a sigma bond (usually C–H or C–C) with an adjacent empty (or partially filled) non-bonding p-orbital or antibonding n orbital or filled π orbital, to give an extended molecular orbital that increases the stability of the system. Only electrons in bonds that are β to the positively charged carbon can stabilize a carbocation by hyperconjugation.

 

Hope this helped you immensely...

All the Very Best & Good Luck to you ...

Regards,

AskIITians Expert,

Godfrey Classic Prince

IIT-Madras

Please approve my answer if you liked it by clicking on "Yes" given below...!!Smile